OpenStudy (anonymous):

An electric scooter has a battery capable of supplying 100 Wh of energy. If friction forces and other losses account for 60.0% of the energy usage, what altitude change can a rider achieve when driving in hilly terrain, if the rider and scooter have a combined weight of 910 N?

OpenStudy (anonymous):

we need to do some unit change

OpenStudy (anonymous):

there is physisc section

OpenStudy (anonymous):

100 Watt Hour= 100 Watt * 2600= 260000 Joules

OpenStudy (anonymous):

Take 40% of that(60% is loss) 260000*.4=104000 Joules

OpenStudy (anonymous):

Work= F times distance 104000=910 * x distance=800/7 meter